# Open-Set Deepfake Detection: A Parameter-Efficient Adaptation Method with Forgery Style Mixture

### v2

@@ IV-F2 Grad-CAM map visualization
We visualize the Grad-CAM [86] maps to identify attention regions of the developed face forgery detector for input queries, as shown in Fig. 11. These maps illustrate the results of the ViT-B baseline model and OSDFD across six unforeseen Deepfake datasets: CDF, WDF, DFDC, DFDC-P, DFR, and FFIW. These datasets encompass practical environmental scenarios, including lighting conditions, head poses, image quality, and facial occlusion. Compared to the baseline model, our proposed method accurately highlights informative manipulated regions. The baseline model sometimes focuses on irrelevant non-forgery areas, and these ‘non-interesting’ regions can lead to wrong decisions. Conversely, our method effectively identifies regions with significant forgery artifacts, contributing to superior face forgery detection performance. Specifically, our approach consistently demonstrates outstanding gradient-based localization across diverse environmental variables, such as extreme head poses in WDF, abnormal color hues in DFDC, facial occlusion (glasses) in DFDC-P, and extreme facial area ratios in FFIW. These observations further underscore the outstanding generalizability of our method in real-world applications.
TABLE VIII: Comparison with other local-frequency extractor.
@@
IV-F3 Cross-dataset results using different seeds
To mitigate the influence of different seeds, we train our models for seven trials using seeds 0-6. The cross-dataset AUC and EER performances are shown in Fig. 10 (a) and (b), respectively. The results indicate that our approach consistently achieves promising performance regardless of the different seeds, demonstrating its robustness across varied initializations. This robustness further underscores the superiority of our method from an additional perspective.
@@ V Conclusions and Future Work
This paper introduces a parameter-efficient and generalized approach towards open-set face forgery detection. Our method, OSDFD, incorporates lightweight LoRA and Adapter layers into ViT backbones, updating only these modules during training to learn global and local forgery-aware features efficiently. On the other hand, we propose a novel forgery-style mixture module that augments the forgery source domains. In this vein, our method achieves SOTA generalizability and robustness in face forgery detection. An ablation study further underscores that the designed network modules and objective functions contribute to the final detection performance. Our proposed method, OSDFD, can be flexibly adapted to other backbones in a plug-and-play manner. Further ablation experiments demonstrate that the adopted parameter-efficient training strategy is effective in face forgery detection. Our proposed method will shed light on developing future face forgery detectors for real-world applications.

### v3

@@ IV-G2 Grad-CAM map visualization
We visualize the Grad-CAM [106] maps to identify attention regions of the developed face forgery detector for input queries, as shown in Fig. 10. These maps illustrate the results of the ViT-B baseline model and OSDFD across six unforeseen Deepfake datasets: CDF, WDF, DFDC, DFDC-P, DFR, and FFIW. These datasets encompass practical environmental scenarios, including lighting conditions, head poses, image quality, and facial occlusion. Compared to the baseline model, our proposed method accurately highlights informative manipulated regions. The baseline model sometimes focuses on irrelevant non-forgery areas, and these ‘non-interesting’ regions can lead to wrong decisions. Conversely, our method effectively identifies regions with significant forgery artifacts, contributing to superior face forgery detection performance. Specifically, our approach demonstrates an outstanding gradient-based localization across diverse environmental variables, such as extreme head poses in WDF, abnormal color hues in DFDC, facial occlusion (glasses) in DFDC-P, and extreme facial area ratios in FFIW. These observations further underscore the outstanding generalizability of our method in real-world applications.
TABLE XI: Comparison with other local-frequency extractor.
@@
TABLE XII: Ablation experiment on $\lambda$ (cross-dataset evaluation).
$\lambda$ CDF DFR WDF
AUC EER AUC EER AUC EER
0 79.85 28.86 93.62 14.04 76.82 31.71
0.6 81.33 26.98 94.88 12.69 78.21 30.81
0.8 82.20 25.43 95.28 11.66 79.03 28.31
1.0 83.35 24.47 95.54 9.90 78.53 28.46
1.2 81.93 25.52 95.02 12.51 78.44 29.02
1.4 80.86 24.41 94.66 11.96 77.97 31.05
TABLE XIII: Impacts of different seeds.
CDF DFDC-P DFR WDF FFIW
[…]
IV-G3 Cross-dataset results using different seeds
To mitigate the influence of different seeds, we train our models for seven trials using seeds 0-6. The cross-dataset AUC and EER performances are shown in Table. XIII. The results indicate that our approach achieves promising performance regardless of the different seeds, demonstrating its robustness across varied initializations. This robustness further underscores the superiority of our method from an additional perspective.
@@ V Conclusions and Future Work
This paper introduces a parameter-efficient and generalized approach towards open-set face forgery detection. Our method, OSDFD, incorporates lightweight LoRA and Adapter layers into ViT backbones, updating only these modules during training to learn global and local forgery-aware features efficiently. On the other hand, we propose a novel forgery-style-mixture module that augments the forgery source domains. In this vein, our method achieves SOTA generalizability and robustness in face forgery detection. An ablation study further underscores that the designed network modules and objective functions contribute to the final detection performance. Our proposed method, OSDFD, can be flexibly adapted to other transformer-based backbones in a plug-and-play manner. Further ablation experiments demonstrate that the adopted parameter-efficient training strategy is effective in face forgery detection. Our proposed method will shed light on developing future face forgery detectors for real-world applications.
@@
VI Acknowledgement
This research is supported in part by the National Nature Science Foundation of China (NSFC) under Grant 62502187; in part by the Natural Science Foundation of Jiangxi Province of China under Grant 20252BAC240015; in part by Sichuan Science and Technology Fund 2025ZNSFSC0511; in part by the Research Grant Council (RGC) of Hong Kong SAR, under a GRF Grant 12203124 and an ECS Grant 22201125; in part by the São Paulo Research Foundation (Fapesp) Horus project, Grant #2023/12865-8 and the Brazilian National Council for Scientific and Technological Development (CNPq), Aletheia project, Grant #442229/2024-0. This work was carried out at the Rapid-Rich Object Search (ROSE) Lab, School of Electrical & Electronic Engineering, Nanyang Technological University (NTU), Singapore. This research is supported in part by A*STAR under its OTS Research Programme (Award S24T2TS006). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the A*STAR.